Calculate the cell potential for the following reactions and predict whether the reaction

will take place. If it will take place, write the cell notation.

1. Zn (s) + 2 Ag+


(aq) → Zn+2 (aq) + 2 Ag (s)

2. Cu (s) + Ni+2 (aq) → Cu+2 (aq) + Ni (s)



1
Expert's answer
2021-08-28T06:26:45-0400

The standard electrode potentials for Zinc, silver, copper and nickel ions are -0.763V, 0.800V, 0.337V and -0.230V respectively


1. Electrode potential = 0.763+ 0.800 = 1.563V


2. Electrode potential = 0.230+ 0.337 = 0.567V




Cell notation for 1. Zn/Zn2+ // 2Ag+ / 2Ag


Cell notation for 2. Cu/Cu2+ // Ni2+/Ni